### v1

# Connectivity of partial basis complexes of freely decomposable groups

Abstract.
We show that the complex of partial bases of the free group of rank $n$, where vertices are seen up to conjugation, is Cohen–Macaulay of dimension $n-1$. This positively answers a conjecture raised by Day and Putman. We prove our results in the more general context of freely decomposable groups.
Key words and phrases:
Free groups; Outer automorphisms; Partial bases; Cohen–Macaulay
1. Introduction
Denote by $F_{n}$ the free group of rank $n$. Write $[x]$ for the conjugacy class of an element or subgroup $x$ of $F_{n}$. The complex of partial bases of $F_{n}$ is the complex $\mathcal{B}_{n}$ whose maximal simplices are sets $\{C_{1},\ldots,C_{n}\}$ of conjugacy classes such that there exists a basis $\{v_{1},\ldots,v_{n}\}$ of $F_{n}$ with $C_{i}=[v_{i}]$. It is not hard to show that $\mathcal{B}_{n}$ is a pure simplicial complex of dimension $n-1$.
This complex was studied by Day and Putman in [DP:complexpartialbases]. They showed that $\mathcal{B}_{n}$ is connected for $n\geq 2$, simply connected if $n\geq 3$, and that the orbit complex $\mathcal{B}_{n}/\mathrm{IA}_{n}$ is $(n-2)$-connected, i.e. spherical of dimension $n-1$. Here, $\mathrm{IA}_{n}$ is the Torelli subgroup of $\operatorname{Aut}(F_{n})$, defined as the kernel of the map $\operatorname{Aut}(F_{n})\to\operatorname{GL}_{n}(\mathbb{Z})$. Day–Putman conjectured [DP:complexpartialbases, Conjecture 1.1] that $\mathcal{B}_{n}$ is $(n-2)$-connected as well. The aim of this note is to prove this conjecture:
Theorem 1.1.
The simplicial complex $\mathcal{B}_{n}$ is Cohen–Macaulay of dimension $n-1$.
We prove a more general version of Theorem 1.1, building on work by the first named author and Gupta [BG:Homotopytypecomplex] and generalisations in [Bru:buildingsfreefactor]. In the present text, we will mostly refer to [Bru:buildingsfreefactora], which contains the material of both [Bru:buildingsfreefactor] and [BG:Homotopytypecomplex].
Let $A$ be a finitely generated group that is freely decomposable, i.e. is isomorphic to a non-trivial free product. A free factor system of $A$ is a finite set $\mathcal{A}=\left\{[A_{1}],\ldots,[A_{k}]\right\}$ of conjugacy classes of non-trivial subgroups $A_{i}\leq A$ such that there is a free factorisation $A=H\ast A_{1}\ast\cdots\ast A_{k}$, where $H$ is some free subgroup of $A$. We allow $\mathcal{A}$ to be empty, but note that the empty set is only a free factor system if $A$ is a free group itself. The corank of $\mathcal{A}$ is the rank of a free complement $H$, that is, $\operatorname{crk}(\mathcal{A})=\operatorname{rk}(H)$, which is well-defined by the Grushko Theorem.
[…]
Theorem 1.2.
The poset $\mathcal{FF}(A,\mathcal{A})$ is $c$-connected for
$c=\max\{\ \operatorname{crk}(\mathcal{A})+|\mathcal{A}|-4,\ \operatorname{crk}(\mathcal{A})-2\ \}.$
For $A=F_{n}$ and $\mathcal{A}=\emptyset$, we have $c=n-2$ and the result is [BG:Homotopytypecomplex, Theorem B]. In [Bru:buildingsfreefactora, Proposition 4.30], it was shown that $\mathcal{FF}(A,\mathcal{A})$ is homotopy equivalent to the simplicial boundary of relative versions of Culler–Vogtmann Outer space [CV:Moduligraphsautomorphisms] as defined by Guirardel–Levitt [GL:outerspacefree], so Theorem 1.2 implies that these boundaries are highly connected.
The frame complex associated to $(A,\mathcal{A})$, denoted by $\mathcal{PB}(A,\mathcal{A})$, is the subposet of $\mathcal{FF}(A,\mathcal{A})\cup\{[A]\}$ given by free factor systems $\mathcal{A}^{\prime}$ such that for each $[A^{\prime}]\in\mathcal{A}^{\prime}$, we either have $[A^{\prime}]\in\mathcal{A}$ or $A^{\prime}\cong\mathbb{Z}$ is a free subgroup of $A$ of rank $1$. Note that $\{[A]\}\in\mathcal{PB}(A,\mathcal{A})$ if and only if $\mathcal{A}=\emptyset$ and $A$ is a free group of rank $1$, and in such a case $\mathcal{PB}(A,\mathcal{A})=\{[A]\}$. In general, the elements of $\mathcal{PB}(A,\mathcal{A})$ are of the form $\left\{[\langle v_{1}\rangle],\ldots,[\langle v_{i}\rangle]\right\}\sqcup\mathcal{A}$, where $\{v_{1},\ldots,v_{i}\}$ is a partial basis of the free group $H$ in a decomposition $A=H\ast A_{1}\ast\cdots\ast A_{k}$. The poset relation of $\mathcal{FF}(A,\mathcal{A})$ restricts to the subset relation on $\mathcal{PB}(A,\mathcal{A})$, so we can see $\mathcal{PB}(A,\mathcal{A})$ as a simplicial complex with vertices of the form $\left\{[\langle v\rangle]\right\}\sqcup\mathcal{A}$.
Theorem 1.3.
The complex $\mathcal{PB}(A,\mathcal{A})$ is Cohen–Macaulay of dimension $\operatorname{crk}(\mathcal{A})-1$.
We deduce Theorem 1.3 from Theorem 1.2 by showing that the inclusion map $\mathcal{PB}(A,\mathcal{A})\to\mathcal{FF}(A,\mathcal{A})$ is highly connected. For $A=F_{n}$ and $\mathcal{A}=\emptyset$, we have $\operatorname{crk}(\mathcal{A})=n$ and $\mathcal{PB}(F_{n},\emptyset)$ is an inflation of the complex $\mathcal{B}_{n}$ (see [BWW:PosetFiberTheorems, Section 6]), so Theorem 1.1 is an immediate consequence of Theorem 1.3.
Related work
The complex $\mathcal{B}_{n}$ comes with a natural action of $\operatorname{Out}(F_{n})$, the outer automorphism group of the free group. More generally, relative outer automorphism groups $\operatorname{Out}(A,\mathcal{A}^{t})$ (see [Bru:buildingsfreefactora, Section 4.1.2]) act on $\mathcal{PB}(A,\mathcal{A})$ and $\mathcal{FF}(A,\mathcal{A})$. We would like to mention that variants of the complexes $\mathcal{B}_{n}$ and $\mathcal{PB}(A,\mathcal{A})$ related to several other groups are known to be highly connected as well: An $\operatorname{Aut}(F_{n})$-version of Theorem 1.1 was established by Sadofschi Costa [SC:complexOfPartialBases], who showed that the simplicial complex whose simplices are the partial bases of $F_{n}$ (without modding out by the conjugation action on vertices) is Cohen–Macaulay of dimension $n-1$. Our paper uses different techniques and is not based on his results. Furthermore, Maazen [Maa:Homologystabilitygeneral] showed that a $\operatorname{GL}_{n}$-version of $\mathcal{B}_{n}$ is Cohen–Macaulay; Harer [Har:Stabilityhomologymapping] proved that a version related to the mapping class group of a surface is highly connected (cf. the introduction of [DP:complexpartialbases]); Putman studied a version related to the integral symplectic group $\operatorname{Sp}_{2n}(\mathbb{Z})$ [Brueck2023, Put:infinitepresentationTorelli].
Acknowledgements
The first author was supported by the Deutsche Forschungsgemeinschaft (Project 427320536–SFB 1442) and Germany’s Excellence Strategy grant EXC 2044–390685587. The second author was supported by the Alexander von Humboldt Stiftung, and the FWO grant 12K1223N.
2. Posets of graphs
For the proof of Theorem 1.2 in Section 3, we compare $\mathcal{FF}(A,\mathcal{A})$ with a poset of free splittings of $A$, and study connectivity of fibres. These fibres coincide with certain finite posets of graphs, which we show to be highly connected in the present section.
2.1. Posets and wedges of spheres
We call a map $f\colon X\to Y$ between two posets a poset map if $x\leq y$ implies $f(x)\leq f(y)$. If $Y$ is a subposet of $X$, we call $f\colon X\to Y$ monotone if either for all $x\in X$ we have $f(x)\leq x$, or for all $x\in X$ we have $f(x)\geq x$. By [Qui:Homotopypropertiesposet, 1.3], such a monotone poset map $f$ defines a homotopy equivalence from $X$ to its image $f(X)$.
We consider a point to be a (trivial) wedge of $d$-spheres for all $d\in\mathbb{Z}$, and for $d<0$, we consider a non-contractible wedge of $d$-spheres to be the same as the empty set.
2.2. Graphs
In what follows, by a graph we mean a $1$-dimensional compact CW-complex with no isolated vertices. For a graph $G$, we denote by $V(G)$ its set of vertices, and by $E(G)$ its set of edges. Note that multiple edges and loops are allowed. By a subgraph we mean an edge-induced subgraph. That is, subgraphs of $G$ are in one-to-one correspondence with subsets of $E(G)$. Recall that the valence of a vertex $v\in V(G)$ is the number of connected components obtained by deleting $v$ from a small enough neighbourhood of $v$, i.e. the number of half-edges adjacent at $v$. An edge that has an endpoint with valence one is called a leaf. An edge is said to be separating if removing any point of its interior disconnects the graph. In particular, we consider leaves to be separating. The rank of a connected graph $G$, denoted by $\operatorname{rk}(G)$, is the rank of its fundamental group.
A labelled graph is a pair $(G,l)$ where $G$ is a graph, $k$ is a non-negative integer, and $l:\{1,\ldots,k\}\to V(G)$ is an injective function. We call the image of $l$ the labelled vertices of $G$ and write $k(l)$ for the number of labelled vertices. A core subgraph of a labelled graph $(G,l)$ is a non-empty subgraph such that every vertex of valence one is labelled. Observe that every labelled graph $(H,l)$ contains a unique maximal core subgraph $\mathring{H}$: this is the subgraph obtained by repeatedly removing leaves with non-labelled endpoints.
Let $e\in E(G)$ be an edge of a labelled graph $(G,l)$. We write $G\setminus e$ for the subgraph of $G$ on $E(G)\setminus\{e\}$ and $G/e$ for the graph obtained from $G$ by collapsing $e$. Both $G\setminus e$ and $G/e$ have obvious labellings induced by $l$. For example, if $e$ is a leaf with a labelled endpoint of valence one and an unlabelled endpoint of higher valence, then $G/e$ and $G\setminus e$ are homeomorphic, but $G/e$ has one more labelled vertex than $G\setminus e$ (namely the one coming from the higher valence endpoint of $e$ in $G$). See Figure 1.
$G$$e$$G/e$$G\setminus e$
Figure 1. A graph $(G,l)$ with labelled vertices in blue (left). It has a leaf $e$ whose valence-one vertex is labelled. The graphs $G/e$ (middle), and $G\setminus e$ (right) are homeomorphic but do not have the same induced labelling.
2.3. Posets of core graphs
Let $\operatorname{C}(G,l)$ be the poset of proper core subgraphs of a labelled graph $(G,l)$, ordered by inclusion. Let $\operatorname{X}(G,l)$ be the poset of proper subgraphs of $G$ such that at least one connected component has non-trivial fundamental group or contains at least two labelled vertices. The poset $\operatorname{C}(G,l)$ is contained in $\operatorname{X}(G,l)$, and indeed $\operatorname{X}(G,l)$ is the upward-closure of $\operatorname{C}(G,l)$ in the poset of proper subgraphs of $G$.
Lemma 2.1.
Let $(G,l)$ be a labelled graph. Then $\operatorname{X}(G,l)$ deformation retracts to $\operatorname{C}(G,l)$.
Proof.
This is because every $H\in\operatorname{X}(G,l)$ contains a unique maximal core subgraph $\mathring{H}\in\operatorname{C}(G,l)$. The assignment $H\mapsto\mathring{H}$ defines a monotone poset map $\operatorname{X}(G,l)\to\operatorname{C}(G,l)$ such that $H\geq\mathring{H}$ and restricts to the identity on $\operatorname{C}(G,l)$. Hence $\operatorname{X}(G,l)$ deformation retracts to $\operatorname{C}(G,l)$ (see [Bru:buildingsfreefactora, Corollary 2.5]). ∎
In [Bru:buildingsfreefactora, Proposition 4.14], it was shown that for $k(l)=0$, the poset $\operatorname{X}(G,l)$ is contractible if and only if $G$ has a separating edge. The following definition is set up in such a way that this result generalises to the cases $k(l)>0$.
Definition 2.2.
We say that an edge $e$ of $(G,l)$ is $l$-separating if it satisfies the following three conditions:
  1. (1)
    The edge $e$ is separating.
  2. (2)
    One of the components of $G\setminus e^{\circ}$ contains all the labelled vertices.
  3. (3)
    The graph $G\setminus e$ has at least one connected component that has non-trivial fundamental group or contains at least two labelled vertices, i.e. $G\setminus e\in\operatorname{X}(G,l)$.
Note that for $k\leq 1$, an edge is $l$-separating if and only if it is separating and $G$ has non-trivial fundamental group. In particular, the $l$-separating edges of $(G,l)$ are the same as the $\emptyset$-separating edges of the trivially labelled graph $(G,\emptyset)$.
Lemma 2.3.
Let $(G,l)$ be a connected labelled graph with $k=k(l)$ labelled points and $\operatorname{rk}(G)=n$. Let
$d(G,l)=\begin{cases}n+k-3&\text{if }k\geq 1,\\ n-2&\text{if }k=0.\end{cases}$
Then $\operatorname{X}(G,l)$ is homotopy equivalent to a wedge of $d(G,l)$-spheres. It is contractible if and only if there is an $l$-separating edge.
To explain the case distinction, note that if $k(l)=1$, then $\operatorname{X}(G,l)=\operatorname{X}(G,\emptyset)$. In [Bru:buildingsfreefactora, Proposition 4.14], Lemma 2.3 was deduced from [Vog:Localstructuresome, Proposition 2.2] for the case $k=0$, $n\geq 2$ under the assumption that every vertex of $G$ has valence at least three.
Proof of Lemma 2.3
Suppose first that $G$ contains an $l$-separating edge $e$. Then by Item 3 of Definition 2.2, $G\setminus e$ is an element of $\operatorname{X}(G,l)$. Item 2 of Definition 2.2 implies that for all $H\in\operatorname{X}(G,l)$, also $H\setminus\left\{e\right\}$ is an element of $\operatorname{X}(G,l)$. The assignment $H\mapsto H\setminus\left\{e\right\}$ gives a monotone poset map $\operatorname{X}(G,l)\to\operatorname{X}(G,l)$, so a homotopy equivalence to its image. The image has a maximal element given by $G\setminus\left\{e\right\}$, so it is contractible. Hence, also $\operatorname{X}(G,l)$ is contractible.
Now assume that $G$ contains no $l$-separating edges. We prove that $\operatorname{X}(G,l)$ is a non-trivial wedge of $d(G,l)$-spheres by induction on the number of edges of $G$. If $G$ has just a single edge, then this edge is not $l$-separating and $\operatorname{X}(G,l)$ is empty, so in particular non-contractible. (This is true independently of whether the edge forms a loop and whether it has labelled endpoints or not.) Hence, the claim is true. In what follows, we will assume that there are at least two edges.
Removing loops
Suppose that $G$ contains a loop $e$. Then the assumption that $G$ has at least two edges and that none of them is $l$-separating implies that $G\setminus e$ is an element of $\operatorname{X}(G,l)$. The poset $\operatorname{X}(G,l)$ can be written as the union of two subposets. Let
$X_{1}=\operatorname{X}(G\setminus e,l^{\prime})\cup\left\{G\setminus e\right\},$
where $l^{\prime}$ is the labelling induced by $l$. This subposet has $G\setminus e$ as its unique maximal element and hence is contractible. The second subposet is given by $\operatorname{X}(G\setminus e,l^{\prime})$ together with all proper subgraphs of $G$ containing $e$:
$X_{2}=\operatorname{X}(G\setminus e,l^{\prime})\cup\left\{H\subsetneq G\mid e\in H\right\}.$
This poset is contractible as well: The assignment $H\mapsto H\cup\left\{e\right\}$ gives a monotone poset map $X_{2}\to X_{2}$ and its image is contractible because it has a unique minimal element $\left\{e\right\}$. Both $X_{1}$ and $X_{2}$ are downward closed, so
$|\operatorname{X}(G,l)|=|X_{1}|\cup_{|X_{1}\cap X_{2}|}|X_{2}|,\text{ where }X_{1}\cap X_{2}=\operatorname{X}(G\setminus e,l^{\prime}).$
Thus $\operatorname{X}(G,l)$ is homotopy equivalent to the topological suspension of $\operatorname{X}(G\setminus e,l^{\prime})$. As $G$ has no $l$-separating edge, also $G\setminus e$ has no $l^{\prime}$-separating edge. The graph $G\setminus e$ is connected, so by induction, $\operatorname{X}(G\setminus e,l^{\prime})$ is homotopy equivalent to a non-trivial wedge of $d(G\setminus e,l^{\prime})$-spheres. Furthermore, $G\setminus e$ has rank one less than $G$ but the same number of labelled points, i.e. $\operatorname{rk}(G\setminus e)=n-1$ and $k(l^{\prime})=k$. Hence, $d(G\setminus e,l^{\prime})=d(G,l)-1$. This implies the claim for $\operatorname{X}(G,l)$.
Loop-free graphs
By the previous paragraph, we now can assume that $G$ contains no loops. Take an edge $e$ which is either non-separating (and not a loop) or a leaf. Let $Y_{1}:=\operatorname{X}(G,l)\setminus\left\{G\setminus e\right\}$ and denote by $(G\setminus e,l^{\prime})$ the labelled graph obtained from $G$ by removing $e$.
Claim 1.
The poset $Y_{1}$ is homotopy equivalent to a wedge of $d(G,l)$-spheres. If $Y_{1}$ is contractible, then $Y_{1}\neq\operatorname{X}(G,l)$ and $G\setminus e$ has no $l^{\prime}$-separating edge.
Proof.
To prove this, we consider two cases. The first is that $e$ has both endpoints labelled. As $G$ has more than one edge, this implies that $\left\{e\right\}\in Y_{1}$. We get a monotone map $Y_{1}\to Y_{1}$ by the assignment $H\mapsto H\cup\left\{e\right\}$. Its image has a unique minimal element $\left\{e\right\}$, so $Y_{1}$ is contractible. Furthermore, as $G$ has no $l$-separating edge, the graph $G\setminus e$ is an element of $\operatorname{X}(G,l)$, i.e. $Y_{1}\neq\operatorname{X}(G,l)$, and $G\setminus e$ has no $l^{\prime}$-separating edge.
The second case is that at least one of the endpoints of $e$ is unlabelled. In this case, for all $H\in Y_{1}$, the image of $H$ in $G/e$ is also an element of $\operatorname{X}(G/e,\tilde{l})$, where $\tilde{l}$ is the labelling induced by $l$. This gives a natural poset map
$f\colon Y_{1}$ $\to\operatorname{X}(G/e,\tilde{l})$
that sends a subgraph to its image in the quotient $G/e$. We define a poset map in the other direction as follows: Let $v_{e}$ be the vertex of $G/e$ to which $e$ was collapsed. Then define
$g\colon\operatorname{X}(G/e,\tilde{l})$ $\to Y_{1}$
$K$$\mapsto\begin{cases}K\cup\left\{e\right\}&v_{e}\in K,\\ K&\text{otherwise}.\end{cases}$
Here we have identified $K$ with the set of corresponding edges in $G$. We have $g\circ f(H)\supseteq H$ and $f\circ g(K)=K$. Hence, $Y_{1}$ and $\operatorname{X}(G/e,\tilde{l})$ have the same homotopy type (see Section 2.1). The graph $G/e$ has no $\tilde{l}$-separating edges (since $(G,l)$ has no $l$-separating edge), so by induction, we see that $Y_{1}\simeq\operatorname{X}(G/e,\tilde{l})$ is homotopy equivalent to a non-trivial wedge of $d(G/e,\tilde{l})$-spheres. Since $e$ is not a loop, we have $\operatorname{rk}(G/e)=\operatorname{rk}(G)$. We also have $k(\tilde{l})=k(l)$ because not both endpoints of $e$ are labelled. Hence, $d(G/e,\tilde{l})=d(G,l)$. ∎
Now we use this subposet to show the desired connectivity for $\operatorname{X}(G,l)$. If $Y_{1}=\operatorname{X}(G,l)$ then the result immediately follows from Claim 1. Thus we may assume that $G\setminus e\in\operatorname{X}(G,l)$. It follows that $\operatorname{X}(G,l)$ is obtained from $Y_{1}=\operatorname{X}(G,l)\setminus\left\{G\setminus e\right\}$ by attaching the star of $G\setminus e$ along $\operatorname{lk}(G\setminus e)$. That is, if $Y_{2}$ is the downward closure of $G\setminus e$ in $\operatorname{X}(G,l)$, then
$|\operatorname{X}(G,l)|=|Y_{1}|\cup_{|Y_{1}\cap Y_{2}|}|Y_{2}|,\text{ where }Y_{1}\cap Y_{2}=\operatorname{lk}(G\setminus e).$
The poset $Y_{2}$ is contractible as it has the unique maximal element $G\setminus e$, and $Y_{1}$ is homotopy equivalent to a wedge of $d(G,l)$-spheres by Claim 1. Hence, the following claim implies that $\operatorname{X}(G,l)$ is homotopy equivalent to a non-trivial wedge of $d(G,l)$-spheres and finishes the proof of Lemma 2.3.
Claim 2.
The poset $\operatorname{lk}(G\setminus e)$ is homotopy equivalent to a wedge of $(d(G,l)-1)$-spheres and either $\operatorname{lk}(G\setminus e)$ or $Y_{1}$ are non-contractible.
Proof.
The link $\operatorname{lk}(G\setminus e)$ consists of all $H\in\operatorname{X}(G,l)$ such that $H<G\setminus e$. It is isomorphic to $\operatorname{X}(G\setminus e,l^{\prime})$, where $l^{\prime}$ is the labelling induced by $l$. As $e$ is either non-separating or a leaf, $G\setminus e$ is connected. Hence, by induction, $\operatorname{X}(G\setminus e,l^{\prime})\cong\operatorname{lk}(G\setminus e)$ is homotopy equivalent to a wedge of $d(G\setminus e,l^{\prime})$-spheres and non-contractible if $G\setminus e$ contains no $l^{\prime}$-separating edge. By Claim 1, either this is the case or $Y_{1}$ is non-contractible. It remains to show that either $d(G\setminus e,l^{\prime})$ and $d(G,l)-1$ are equal or both are smaller than 00.
If $e$ is non-separating, then $\operatorname{rk}(G\setminus e)=\operatorname{rk}(G)-1$ and $k(l^{\prime})=k$. Hence the equality $d(G\setminus e,l^{\prime})=d(G,l)-1$ holds. Now assume that $e$ is a leaf. If $k\leq 1$ and $n=0$, then $\operatorname{X}(G\setminus e,l^{\prime})=\emptyset$. This is consistent with $d(G,l)=-2$. Otherwise, $(G,l)$ not having $l$-separating edges implies that the valence-one vertex of $e$ is labelled and that $k\geq 2$. Hence, we have $k(l^{\prime})=k-1\geq 1$ and $\operatorname{rk}(G\setminus e)=\operatorname{rk}(G)$, so we again get equality $d(G\setminus e,l^{\prime})=d(G,l)-1$. ∎
3. Proof of the main theorems
In this section, we prove Theorem 1.1, Theorem 1.3 and Theorem 1.2. We will frequently use Quillen’s fibre theorem as stated in [PW:Decompositions, Theorem 5.3] to show that certain maps are highly connected.
Throughout the section, we fix a finitely generated group $A$ and a (possibly empty) free factor system $\mathcal{A}$ of $A$. We set
$n=\operatorname{crk}(\mathcal{A})\text{ and }k=|\mathcal{A}|.$
We denote by $\mathcal{FS}=\mathcal{FS}(A,\mathcal{A})$ the poset of free splittings of $A$ relative to $\mathcal{A}$. Its elements can be seen as pairs $S=(\mathbb{G},\mathfrak{m})$, where $\mathbb{G}$ is a graph of groups with underlying labelled graph a connected core graph in the sense of Section 2, and where $\mathfrak{m}\colon A\to\pi_{1}(\mathbb{G})$ is an isomorphism which is well-defined up to inner automorphisms. The poset relation is given by saying that $S>S^{\prime}$ if $S^{\prime}$ is obtained from $S$ by collapsing a subgraph. For $S\in\mathcal{FS}$, we denote by $\mathcal{V}(S)$ the free factor system given by the conjugacy classes of its vertex groups. By definition, we have $\mathcal{V}(S)\in\mathcal{FF}\cup\{\mathcal{A}\}$, i.e. $\mathcal{A}\sqsubseteq\mathcal{V}(S)$. We refer the reader to [Bru:buildingsfreefactora, Section 4.1] for more detailed definitions.
We set $L=L(A,\mathcal{A})=\{S\in\mathcal{FF}\mathrel{{\>:\>}}\mathcal{V}(S)=\mathcal{A}\}$ and $b\mathcal{FS}=\mathcal{FS}\setminus L$. Here, $L$ is the spine of relative Outer space $\mathcal{O}(A,\mathcal{A})$ in the sense of [GL:outerspacefree], and $b\mathcal{FS}$ is its simplicial boundary. This simplicial boundary has the homotopy type of the poset $\mathcal{FF}$ that we are interested in:
Lemma 3.1.
We have a homotopy equivalence $\mathcal{V}:b\mathcal{FS}\to\mathcal{FF}^{\operatorname{op}}$.
Proof.
This is [Bru:buildingsfreefactora, Proposition 4.30.1]. ∎
Let $Z\subseteq L\times b\mathcal{FS}$ be the subposet consisting of all pairs $(S,S^{\prime})$ where $S^{\prime}$ is obtained from $S$ by collapsing a proper core subgraph. We have projections $p_{1}\colon Z\to L$ and $p_{2}\colon Z\to b\mathcal{FS}$.
Lemma 3.2.
The projection $p_{2}:Z\to b\mathcal{FS}$ is a homotopy equivalence.
Proof.
This is part of [Bru:buildingsfreefactora, Corollary 4.35]. ∎
Now we prove that $p_{1}$ is highly connected. Section 4.6.4 of [Bru:buildingsfreefactora] studies this map for the case $A=F_{n}$ and $\mathcal{A}=\emptyset$. Here we use Lemma 2.3 to generalise this to an arbitrary pair $(A,\mathcal{A})$.
Lemma 3.3.
The projection $p_{1}:Z\to L$ is an $e(n,k)$-equivalence, where
$e(n,k)=\begin{cases}n+k-3&\text{if }k\geq 2,\\ n-1&\text{if }k=1,\\ n-2&\text{if }k=0.\end{cases}$
Note that $e(n,k)$ agrees with $d(G,l)$ from Lemma 2.3 except if $k=1$. This case distinction will become clear in the following proof.
Proof of Lemma 3.3.
Let $S\in L$ and let $(G,l)$ be the underlying labelled graph. Note that $G$ has exactly $k$ labelled vertices and $\operatorname{rk}(G)=n$. By [Bru:buildingsfreefactora, Lemma 4.36], the fibre $p_{1}^{-1}(L_{\leq S})$ is homotopy equivalent to $p_{1}^{-1}(S)$. (The proof of [Bru:buildingsfreefactora, Lemma 4.36] is formulated for $A=F_{n}$, $\mathcal{A}=\emptyset$, but it generalises verbatim to the setup here.) The latter is exactly the poset $\operatorname{C}(G,l)$ defined in Section 2. This is $(d(G,l)-1)$-connected by Lemma 2.1 and Lemma 2.3. For $k\neq 1$, we have $d(G,l)=e(n,k)$ and the claim directly follows from Quillen’s fibre theorem [PW:Decompositions, Theorem 5.3].
For $k=1$, we have $d(G,l)=n-2$, so the same argument would give us one degree of connectivity less than claimed. Assume we are in this case. We will prove that
$p_{1}^{-1}(L_{\leq S})*L_{>S}$
[…]
$G$$v$$\widetilde{G}$$v$
Figure 2. Left: A graph $G$ with a labelled vertex $v$ (in blue) of valence at least two. Right: The graph $\widetilde{G}$ obtained by pushing away the labelling from $v$ and creating a leaf. We have $G<\widetilde{G}$ and every non-labelled vertex of $\widetilde{G}$ has valence at least three.
As a corollary, we get Theorem 1.2, which states that $\mathcal{FF}$ is $(n+k-4)$-connected for $k\geq 2$ and $(n-2)$-connected for $k\in\left\{0,1\right\}$:
Proof of Theorem 1.2.
As pointed out in the introduction, the case $k=0$, i.e. $\mathcal{A}=\emptyset$, only occurs if $A$ is a free group. In this case, $n=\operatorname{rk}(A)$ and Theorem 1.2 follows from [BG:Homotopytypecomplex, Theorem B]. Hence, we can assume that $k\neq 0$; we can also assume that $n+k\geq 2$ as there is nothing to show otherwise. Under these assumptions, the poset $L$ is contractible by [GL:outerspacefree, Section 4]. Thus by Lemma 3.1, Lemma 3.2 and Lemma 3.3, $\mathcal{FF}$ is $(n+k-4)$-connected for $k\geq 2$, and $(n-2)$-connected if $k=1$. In any case, $\mathcal{FF}$ is always at least $(n-2)$-connected. ∎
The following is a corollary of [HandelMosher, Lemma 2.5].
Proposition 3.4.
Let $\mathcal{A}^{\prime}\in\mathcal{FF}=\mathcal{FF}(A,\mathcal{A})$. Then
$\mathcal{FF}_{\sqsubseteq\mathcal{A}^{\prime}}\cong\left(\ \prod_{[B]\in\mathcal{A}^{\prime}}\left(\mathcal{FF}(B,\mathcal{A}_{[B]})\cup\{\,\{[B]\},\emptyset\,\}\right)\ \right)\setminus\{\,(\emptyset,\ldots,\emptyset)\,\},$
where $\mathcal{A}_{[B]}=\{\,[C]\in\mathcal{A}\mathrel{{\>:\>}}\{[C]\}\sqsubseteq\{[B]\}\,\}$.
If $\{[B]\}\in\mathcal{FF}(A,\mathcal{A})\cup\{[A]\}$, we write $\operatorname{crk}_{B}(\mathcal{A})$ for the corank of $\mathcal{A}$ viewed as a free factor system of $B$. For $\mathcal{A}^{\prime}\in\mathcal{FF}$, we have
(1) $\operatorname{crk}_{A}(\mathcal{A})=\operatorname{crk}_{A}(\mathcal{A}^{\prime})+\sum_{[B]\in\mathcal{A}^{\prime}}\operatorname{crk}_{B}(\mathcal{A}_{[B]}).$
We are now ready to prove Theorem 1.3, i.e. to show that $\mathcal{PB}$ is Cohen–Macaulay of dimension $n-1$.
Proof of Theorem 1.3.
We argue by induction on $n+k$. If $k=0$, i.e. $\mathcal{A}=\emptyset$, the base case is $n=1$, and here by definition $\mathcal{PB}=\{[A]\}$, which is contractible. Hence we can assume that $n\geq 2$ if $\mathcal{A}=\emptyset$. Furthermore, if $n=0$ then $\mathcal{PB}$ is empty, so we can also assume that $n\geq 1$ if $k\neq 0$. In particular, we always have a poset inclusion $i:\mathcal{PB}\hookrightarrow\mathcal{FF}$.
In view of the previous paragraph, it is enough to show that the inclusion map $i$ is an $(n-1)$-equivalence, and that the codomain is $(n-2)$-connected. To that end, we investigate the fibres and apply Quillen’s fibre theorem.
Let $\mathcal{A}^{\prime}\in\mathcal{FF}\setminus\mathcal{PB}$. By Proposition 3.4, we have
$i^{-1}\big{(}\ \mathcal{FF}_{\sqsubseteq\mathcal{A}^{\prime}}\ \big{)}\cong\left(\ \prod_{[B]\in\mathcal{A}^{\prime}}(\mathcal{PB}(B,\mathcal{A}_{[B]})\cup\{\emptyset\})\ \right)\setminus\{(\emptyset,\ldots,\emptyset)\}.$
This poset is homotopy equivalent to the join
(2) $\bigast_{[B]\in\mathcal{A}^{\prime}}\mathcal{PB}(B,\mathcal{A}_{[B]}).$
Since a class $[B]\in\mathcal{A}^{\prime}$ gives a proper subgroup $B<A$, we have
$\operatorname{crk}_{B}(\mathcal{A}_{[B]})+|\mathcal{A}_{[B]}|<\operatorname{crk}_{A}(\mathcal{A})+|\mathcal{A}|=n+k$
by Equation 1. Therefore, by induction, each $\mathcal{PB}(B,\mathcal{A}_{[B]})$ is $(\operatorname{crk}_{B}(\mathcal{A}_{[B]})-2)$-connected, and so the join of Equation 2 is
$(2|\mathcal{A}^{\prime}|-2)+\sum_{[B]\in\mathcal{A}^{\prime}}(\operatorname{crk}_{B}(\mathcal{A}_{[B]})-2)=n-\operatorname{crk}_{A}(\mathcal{A}^{\prime})-2$
connected.
On the other hand, since $\mathcal{A}\sqsubset\mathcal{A}^{\prime}$, by Theorem 1.2 we have that
$\mathcal{FF}_{\sqsupset\mathcal{A}^{\prime}}=\mathcal{FF}(A,\mathcal{A}^{\prime})$
is at least $(\operatorname{crk}(\mathcal{A}^{\prime})-2)$-connected. Hence the join
$i^{-1}\big{(}\ \mathcal{FF}_{\sqsubseteq\mathcal{A}^{\prime}}\ \big{)}*\mathcal{FF}_{\sqsupset\mathcal{A}^{\prime}}$
is at least $(n-2)$-connected. By Quillen’s fibre theorem, the map $i$ is an $(n-1)$-equivalence.
Finally, the codomain $\mathcal{FF}$ is at least $(n-2)$-connected by Theorem 1.2. Therefore $\mathcal{PB}$ is $(n-2)$-connected.
To conclude with the proof, we show the Cohen–Macaulayness. As noted in the paragraph before Theorem 1.3, we can regard $\mathcal{PB}$ as a simplicial complex. The link of a simplex $\mathcal{A}^{\prime}\in\mathcal{PB}\cup\{\emptyset\}$ can be identified with $\mathcal{PB}(A,\mathcal{A}^{\prime})$, which we saw is $(\operatorname{crk}(\mathcal{A}^{\prime})-2)$-connected. But this link has dimension $\operatorname{crk}(\mathcal{A}^{\prime})-1$, so it is spherical of this dimension. Thus $\mathcal{PB}$ is Cohen–Macaulay of dimension $n-1$. ∎
Proof of Theorem 1.1.
Let $A$ be a free group of finite rank $n$. Since $\mathcal{B}(A)$ is the inflation of $\mathcal{PB}(A,\emptyset)$, and the latter is Cohen-Macaulay of dimension $n-1$ by Theorem 1.3, it follows that $\mathcal{B}(A)$ is also Cohen–Macaulay of dimension $n-1$ (cf. Corollary 6.4 of [BWW:PosetFiberTheorems]). ∎
[…]